# Practical trainable temporal post-processor for multi-state quantum measurement

### v2

@@ D.3.1 Obtaining the linear system for filters
$\mathbf{C}\vec{S}^{(c)}$ $=\begin{pmatrix}\sum_{c^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}(\bm{\vec{s}}^{(c^{\prime})})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}\!+\!{\color[rgb]{0.0,0.0,0.0}\bar{\mathbf{I}}}&\sum_{c^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}(\bm{\vec{s}}^{(c^{\prime})})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}&C\end{pmatrix}\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ 1\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}[(\bm{\vec{s}}^{(c^{\prime})})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}]+{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}+\sum_{c^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}[(\bm{\vec{s}}^{(c^{\prime})})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}]+C\end{pmatrix}$ (88)
@@
$O_{cc^{\prime}}=(\bm{\vec{s}}^{(c^{\prime})})^{T}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}{\color[rgb]{0.0,0.0,0.0}=(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{V}^{-1}\bm{\vec{s}}^{(c)}},$ (89)
@@
$\mathbf{C}\vec{S}^{(c)}$ $=\begin{pmatrix}\sum_{c^{\prime}}O_{cc^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}+\sum_{c^{\prime}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}+{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c)}\\ \sum_{c^{\prime}}O_{cc^{\prime}}+C\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}\left[O_{cc^{\prime}}+1\right]\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\end{pmatrix}-\begin{pmatrix}\bm{\vec{0}}\\ \sum_{c^{\prime}}\delta_{cc^{\prime}}\end{pmatrix}$
$=\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\begin{pmatrix}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(c^{\prime})}\\ 1\end{pmatrix}-\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}$ (90)
@@
$\vec{S}^{(c)}$ $=\sum_{c^{\prime}}(\mathbf{C}-\lambda\mathbf{I})^{-1}(M_{cc^{\prime}}-\lambda\mathbf{I}\delta_{cc^{\prime}})\vec{S}^{(c^{\prime})}-(\mathbf{C}-\lambda\mathbf{I})^{-1}{\color[rgb]{0.0,0.0,0.0}\vec{n}}$
$=\sum_{c^{\prime}}(M_{cc^{\prime}}-\lambda\mathbf{I}\delta_{cc^{\prime}})\vec{F}_{c^{\prime}}-(\mathbf{C}-\lambda\mathbf{I})^{-1}{\color[rgb]{0.0,0.0,0.0}\vec{n}},$ (93)
@@ D.3.2 Solving the linear system for filters
$({\color[rgb]{0.0,0.0,0.0}\mathbf{L}}^{-1})^{T}\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(P_{p})}$ (102)
@@
$\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}^{T}}{\color[rgb]{0.0,0.0,0.0}\mathbf{L}}\bm{\vec{s}}^{(P_{p})}\implies\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}{\color[rgb]{0.0,0.0,0.0}\mathbf{V}}^{-1}\bm{\vec{s}}^{(P_{p})}$ (103)
@@
An alternative, more practical form of the learned filters can be extracted by transition from the representation in terms of difference vectors $\vec{S}^{P_{p}}$, to the individual traces $\vec{S}^{(c)}$, using Eq. (97). We find:
$\bm{\vec{f}}_{c}=\mathbf{Q}^{-1}_{c1}{\color[rgb]{0.0,0.0,0.0}\mathbf{V}}^{-1}\bm{\vec{s}}^{(1)}+\sum_{p=2}^{C-1}\left[\mathbf{Q}^{-1}_{cp}-\mathbf{Q}^{-1}_{c(p-1)}\right]{\color[rgb]{0.0,0.0,0.0}\mathbf{V}}^{-1}\bm{\vec{s}}^{(p)}-\mathbf{Q}^{-1}_{c(C-1)}{\color[rgb]{0.0,0.0,0.0}\mathbf{V}}^{-1}\bm{\vec{s}}^{(C)}$ (105)
@@
$\bm{\vec{f}}_{c}=\sum_{c=1}^{C}C_{cp}{\color[rgb]{0.0,0.0,0.0}\mathbf{V}}^{-1}\bm{\vec{s}}^{(p)},\leavevmode\nobreak\ C_{cp}=\begin{cases}+\mathbf{Q}^{-1}_{c1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm if}\leavevmode\nobreak\ c=1,\\ -\mathbf{Q}^{-1}_{c(C-1)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm if}\leavevmode\nobreak\ c=C,\\ \mathbf{Q}^{-1}_{cp}-\mathbf{Q}^{-1}_{c(p-1)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm otherwise}.\end{cases}$ (106)
@@ D.4 Reduction to standard matched filter for binary classification ($C=2$)
$\begin{pmatrix}\bm{\vec{f}}_{1}\\ \mathbf{b}_{1}\end{pmatrix}=\frac{{\color[rgb]{0.0,0.0,0.0}\mathbf{V}^{-1}}}{M_{11}-M_{21}-(M_{12}-M_{22})}\begin{pmatrix}\bm{\vec{s}}^{(1)}-\bm{\vec{s}}^{(2)}\\ 0\end{pmatrix}-\frac{M_{21}-M_{22}}{M_{11}-M_{21}-(M_{12}-M_{22})}\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}$ (108)
D.5 TPP-learned optimal filters for multi-state classification under Gaussian white noise
@@
D.6 Comparison of TPP against $C-1$ instances of FGDA
@@
D.7 Semi-analytic TPP-learned optimal filters beyond Gaussian white noise conditions
@@
We start with a simple case where the required $\mathbf{V}^{-1}$ can be computed analytically. Consider the case of heterodyne measurement $N_{\rm O}=2$ but where the two measured observables (quadrature time series $\vec{I}$ and $\vec{Q}$) have stationary but distinct variances $\sigma_{I}^{2},\sigma_{Q}^{2}$ respectively. In this case, $\mathbf{V}$ takes the simple form:
@@
where $\tilde{\mathbf{I}}$ is the identity matrix in $\mathbb{R}^{N_{\rm T}}$. Of course, this form of $\mathbf{V}$ can be straightforwardly inverted:
@@
Figure 12: Comparison of semi-analytic and exact TPP filters under general noise correlation conditions. Filters are shown for (a) simulated data where different quadrature time series $\vec{I},\vec{Q}$ have different variances $\sigma^{2}_{I},\sigma^{2}_{Q}$ respectively, and (b) simulated data from a phase-preserving quantum-limited amplifier, as in Sec. V.1 of the main text. Excellent agreement is observed between the exact general TPP filter and the semi-analytic general TPP filter, while both are markedly different from the black curves, which show TPP filters under the assumption of uniform (namely, observable-independent) Gaussian white noise.
@@ E.1 Classification performance versus increasing signal amplitude for real qubit readout
In Fig. 13, we now plot $\mathcal{N}_{\rm FGDA}$ against $\mathcal{N}_{\rm TPP}$ for the two dispersive qubit-cavity systems that were analyzed as a function of measurement tone amplitude in the main text. The datapoint at $\mathcal{N}_{\rm FGDA}=\mathcal{N}_{\rm TPP}=1$ for each dataset corresponds to the lowest amplitude, by construction of Eq. (113). We next note that a few datapoints fall into the category where both $\mathcal{N}_{\rm TPP}>1$ and $\mathcal{N}_{\rm FGDA}>1$, indicating a reduction in fidelity with increasing signal amplitude using either classification scheme. Here, non-idealities such as enhanced qubit transitions degrade the measured data, which neither classification scheme is able to overcome. We note that here we still have $\mathcal{N}_{\rm FGDA}>\mathcal{N}_{\rm TPP}$, so that the FGDA performance is worse than the TPP.
@@
Figure 13: Classification performance of TPP versus FGDA as a function of increasing measurement tone (signal) amplitude for readout of real qubits. Same data as Fig. 3 of the main text, but now plotting $\mathcal{N}$, the classification infidelity normalized by infidelity at lowest signal amplitude for each dispersive qubit-cavity system shown; see Eq. (113. As before, more opaque markers again indicate stronger measurement tone amplitudes.

### v3

@@ D.3.1 Obtaining the linear system for filters
$\mathbf{C}\vec{S}^{(c)}$ $=\begin{pmatrix}\sum_{c^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{L}^{T}\!+\!\bar{\mathbf{I}}&\sum_{c^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{L}^{T}&C\end{pmatrix}\begin{pmatrix}\mathbf{L}\bm{\vec{s}}^{(c)}\\ 1\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}[(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{L}^{T}\mathbf{L}\bm{\vec{s}}^{(c)}]+\mathbf{L}\bm{\vec{s}}^{(c)}+\sum_{c^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}[(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{L}^{T}\mathbf{L}\bm{\vec{s}}^{(c)}]+C\end{pmatrix}$ (88)
@@
$O_{cc^{\prime}}=(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{L}^{T}\mathbf{L}\bm{\vec{s}}^{(c)}=(\bm{\vec{s}}^{(c^{\prime})})^{T}\mathbf{V}^{-1}\bm{\vec{s}}^{(c)},$ (89)
@@
$\mathbf{C}\vec{S}^{(c)}$ $=\begin{pmatrix}\sum_{c^{\prime}}O_{cc^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}+\sum_{c^{\prime}}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}+\mathbf{L}\bm{\vec{s}}^{(c)}\\ \sum_{c^{\prime}}O_{cc^{\prime}}+C\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}\left[O_{cc^{\prime}}+1\right]\end{pmatrix}$
$=\begin{pmatrix}\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}\\ \sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\end{pmatrix}-\begin{pmatrix}\bm{\vec{0}}\\ \sum_{c^{\prime}}\delta_{cc^{\prime}}\end{pmatrix}$
$=\sum_{c^{\prime}}\left[O_{cc^{\prime}}+1+\delta_{cc^{\prime}}\right]\begin{pmatrix}\mathbf{L}\bm{\vec{s}}^{(c^{\prime})}\\ 1\end{pmatrix}-\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}$ (90)
@@
$\vec{S}^{(c)}$ $=\sum_{c^{\prime}}(\mathbf{C}-\lambda\mathbf{I})^{-1}(M_{cc^{\prime}}-\lambda\mathbf{I}\delta_{cc^{\prime}})\vec{S}^{(c^{\prime})}-(\mathbf{C}-\lambda\mathbf{I})^{-1}\vec{n}$
$=\sum_{c^{\prime}}(M_{cc^{\prime}}-\lambda\mathbf{I}\delta_{cc^{\prime}})\vec{F}_{c^{\prime}}-(\mathbf{C}-\lambda\mathbf{I})^{-1}\vec{n},$ (93)
@@ D.3.2 Solving the linear system for filters
$(\mathbf{L}^{-1})^{T}\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}\mathbf{L}\bm{\vec{s}}^{(P_{p})}$ (102)
@@
$\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}\mathbf{L}^{T}\mathbf{L}\bm{\vec{s}}^{(P_{p})}\implies\bm{\vec{f}}_{c}=\sum_{p}\mathbf{Q}^{-1}_{cp}\mathbf{V}^{-1}\bm{\vec{s}}^{(P_{p})}$ (103)
@@
An alternative, more practical form of the learned filters can be extracted by transitioning from the representation in terms of difference vectors $\vec{S}^{P_{p}}$, to the individual traces $\vec{S}^{(c)}$, using Eq. (97). We find:
$\bm{\vec{f}}_{c}=\mathbf{Q}^{-1}_{c1}\mathbf{V}^{-1}\bm{\vec{s}}^{(1)}+\sum_{p=2}^{C-1}\left[\mathbf{Q}^{-1}_{cp}-\mathbf{Q}^{-1}_{c(p-1)}\right]\mathbf{V}^{-1}\bm{\vec{s}}^{(p)}-\mathbf{Q}^{-1}_{c(C-1)}\mathbf{V}^{-1}\bm{\vec{s}}^{(C)}$ (105)
@@
$\bm{\vec{f}}_{c}=\sum_{p=1}^{C}C_{cp}\mathbf{V}^{-1}\bm{\vec{s}}^{(p)},\leavevmode\nobreak\ C_{cp}=\begin{cases}+\mathbf{Q}^{-1}_{c1}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm if}\leavevmode\nobreak\ p=1,\\ -\mathbf{Q}^{-1}_{c(C-1)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm if}\leavevmode\nobreak\ p=C,\\ \mathbf{Q}^{-1}_{cp}-\mathbf{Q}^{-1}_{c(p-1)}\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ &{\rm otherwise}.\end{cases}$ (106)
@@ D.4 Reduction to standard matched filter for binary classification ($C=2$)
$\begin{pmatrix}\bm{\vec{f}}_{1}\\ \mathbf{b}_{1}\end{pmatrix}=\frac{\mathbf{V}^{-1}}{M_{11}-M_{21}-(M_{12}-M_{22})}\begin{pmatrix}\bm{\vec{s}}^{(1)}-\bm{\vec{s}}^{(2)}\\ 0\end{pmatrix}-\frac{M_{21}-M_{22}}{M_{11}-M_{21}-(M_{12}-M_{22})}\begin{pmatrix}\bm{\vec{0}}\\ 1\end{pmatrix}$ (108)
D.5 Example: analytic construction of TPP-learned filters for three-state classification ($C=3$)
We now provide an example of the construction of TPP-learned optimal filters for $C=3$ state classification. To compute these filters using Eq. (106), we simply require knowledge of the matrix $\mathbf{Q}$, whose matrix elements are given by Eq. (99). For $C=3$, $\mathbf{Q}\in\mathbb{R}^{2\times 2}$, and the distinct state pairs $P_{p}$ for $p=1,2$ are given by $P_{1}=[1,2],P_{2}=[2,3]$. Then, $\mathbf{Q}$ takes the form:
$\mathbf{Q}=\begin{pmatrix}M_{11}-M_{21}-(M_{13}-M_{23})&&M_{12}-M_{22}-(M_{13}-M_{23})\\ M_{21}-M_{31}-(M_{23}-M_{33})&&M_{22}-M_{32}-(M_{23}-M_{33})\end{pmatrix}$ (109)
and its inverse can hence be easily computed:
$\mathbf{Q}^{-1}=\frac{1}{\det\mathbf{Q}}\begin{pmatrix}M_{22}-M_{32}-(M_{23}-M_{33})&&(M_{13}-M_{23})-(M_{12}-M_{22})\\ (M_{23}-M_{33})-(M_{21}-M_{31})&&M_{11}-M_{21}-(M_{13}-M_{23})\end{pmatrix}$ (110)
Using Eq. (106), we can therefore write for the non-trivial TPP-learned filters:
$\bm{\vec{f}}_{1}$ $=\frac{\mathbf{V}^{-1}}{\det\mathbf{Q}}\left\{\left[M_{22}-M_{32}-(M_{23}-M_{33})\right]\bm{\vec{s}}^{(1)}+\left[M_{13}-M_{12}-(M_{33}-M_{32})\right]\bm{\vec{s}}^{(2)}+\left[M_{12}-M_{22}-(M_{13}-M_{23})\right]\bm{\vec{s}}^{(3)}\right\}$ (111a)
$\bm{\vec{f}}_{2}$ $=\frac{\mathbf{V}^{-1}}{\det\mathbf{Q}}\left\{\left[M_{23}-M_{33}-(M_{21}-M_{31})\right]\bm{\vec{s}}^{(1)}+\left[M_{11}-M_{13}-(M_{31}-M_{33})\right]\bm{\vec{s}}^{(2)}+\left[M_{13}-M_{23}-(M_{11}-M_{21})\right]\bm{\vec{s}}^{(3)}\right\}$ (111b)
Note that the final filter $\bm{\vec{f}}_{3}$ must be defined by the constraint Eq. (16) of the main text (or equivalently Eq. (82) of this appendix section).
D.6 TPP-learned optimal filters for multi-state classification under Gaussian white noise
@@
D.7 Comparison of TPP against $C-1$ instances of FGDA
@@
D.8 Semi-analytic TPP-learned optimal filters beyond Gaussian white noise conditions
@@
We start with a simple case where the required $\mathbf{V}^{-1}$ can be computed analytically. Consider the case of heterodyne measurement $N_{\rm O}=2$ but where the two measured observables (quadrature time series $\vec{I}$ and $\vec{Q}$) have stationary but distinct variances $\sigma_{I}^{2},\sigma_{Q}^{2}$ respectively; for concreteness we assume $\sigma_{Q}^{2}>\sigma_{I}^{2}$. In this case, $\mathbf{V}$ takes the simple form:
@@
where $\tilde{\mathbf{I}}$ is the identity matrix in $\mathbb{R}^{N_{\rm T}\ timesN_{\rm T}}$. Of course, this form of $\mathbf{V}$ can be straightforwardly inverted:
@@
Figure 12: Comparison of semi-analytic and general TPP filters under general noise correlation conditions. Filters are shown for (a) simulated data where different quadrature time series $\vec{I},\vec{Q}$ have different variances $\sigma^{2}_{I},\sigma^{2}_{Q}$ respectively, and (b) simulated data from a phase-preserving quantum-limited amplifier, as in Sec. V.1 of the main text. Excellent agreement is observed between the exact general TPP filter and the semi-analytic general TPP filter, while both are markedly different from the black curves, which show TPP filters under the assumption of uniform (namely, observable-independent) Gaussian white noise.
@@ E.1 Classification performance versus increasing signal amplitude for real qubit readout
In Fig. 13, we now plot $\mathcal{N}_{\rm FGDA}$ against $\mathcal{N}_{\rm TPP}$ for the two dispersive qubit-cavity systems that were analyzed as a function of measurement tone amplitude in the main text. The datapoint at $\mathcal{N}_{\rm FGDA}=\mathcal{N}_{\rm TPP}=1$ for each dataset corresponds to the lowest amplitude, by construction of Eq. (116). We next note that a few datapoints fall into the category where both $\mathcal{N}_{\rm TPP}>1$ and $\mathcal{N}_{\rm FGDA}>1$, indicating increased infidelity with increasing signal amplitude using either classification scheme. Here, non-idealities such as enhanced qubit transitions degrade the measured data, which neither classification scheme is able to overcome. We note that here we still have $\mathcal{N}_{\rm FGDA}>\mathcal{N}_{\rm TPP}$, so that the FGDA performance is worse than the TPP.
@@
Figure 13: Classification performance of TPP versus FGDA as a function of increasing measurement tone (signal) amplitude for readout of real qubits. Same data as Fig. 3 of the main text, but now plotting $\mathcal{N}$, the classification infidelity normalized by infidelity at lowest signal amplitude for each dispersive qubit-cavity system shown; see Eq. (116). As before, more opaque markers again indicate stronger measurement tone amplitudes.